Donna earns twice as much money per month as Omar does. Omar earns $200 more than Alex. Together, the three workers earn $3,320 per month. How much does Omar earn per month?
Answer by macston(5194)

.
A=amount Alex earns; M=amount Omar earns=A+$200; D=amount Donna earns=2M=2A+$400
.
A+M+D=$3320
A+(A+$200)+(2A+$400)=$3320
4A=$2720
A=$680 Alex earns $680 per month.
M=A+$200=$680+$200=$880 ANSWER: Omar earns $880 per month.
.
CHECK:
D=2M=2($880)=$1760
A+M+D=$3320
$680+$880+$1760=$3320
$3320=$3320